If(2a+2)^2+(4b-3)^2+(4c-5)^2=0 then find the value of a^3+b^3+c^3-3abc/a^2+b^2+c^2
Please give me the answer in step wise

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
Sumnof all the squares is 0 only when numbers in brackets zero.
 
Hence
 
2a +2 = 0
a = -1
 
4b -3 = 0
b = ¾
 
4c -5 = 0
c = 5/4
 
Now put these values  in the expression